OpenStudy (anonymous):

Find the quotient. n^ 7 ÷ n ^3

OpenStudy (anonymous):

subtract the exponents

OpenStudy (anonymous):

\[\frac{n^7}{n^3}=n^{7-3}=n^4\]

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

yw